OpenStudy (anonymous):

If y varies inversely as the square of x and y=5 when x=2.5, what is the value of y when x=9?

OpenStudy (whpalmer4):

If \(y\) varies inversely with the square of \(x\), that means we can write \[y = \frac{k}{x^2}\] or \[k=x^2y\]

OpenStudy (whpalmer4):

From that, you should have no trouble finding the value of \(k\) and finding the desired value of \(y\) when \(x=9\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i got 4.5 as my answer

OpenStudy (anonymous):

y=4.5

OpenStudy (whpalmer4):

hmm...not what I get. Want to show your work?

OpenStudy (whpalmer4):

\[k = x^2y\]\[k = (2.5)^2* 5 = 6.25*5 = 31.25\] So the variation equation is \[31.25 = x^2y\]or\[y=\frac{31.25}{x^2}\]What do you get for \(y\) when you plug in \(x=9\) in that equation?

OpenStudy (anonymous):

i was walking with my dad but thank you i was meaning to ask you the next step after plugging in x and y but now i got it

OpenStudy (whpalmer4):

What do you get for your new, improved answer?

OpenStudy (anonymous):

9^2 is 81 \[31.25\div81\] = 0.39

OpenStudy (anonymous):

i rounded it

OpenStudy (whpalmer4):

Yes, much better.

OpenStudy (anonymous):

thank you for your help

OpenStudy (whpalmer4):

you're welcome!